Determine the standard error if a population standard deviation is 31 and sample

size is 19. Round the answer to the two decimal place.


1
Expert's answer
2022-05-05T06:46:36-0400
"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{31}{\sqrt{19}}\approx7.11"